# DSP: Dynamic Sequence Parallelism for Multi-Dimensional Transformers

### v2

@@ 4 Theoretical Analysis
We choose 2D-Transformer as described in Equation 5 as our base model, which is widely employed in real-world applications. To be specific, we use the OpenSora [50] variant of 2D-Transformer, an open-source video generation model, where there are two transformer blocks for two sequence dimensions separately. More details can be found in Appendix A.1. We choose DeepSpeed-Ulysses [20], Megatron-SP [22] and RingAttention [25] as baselines, which represent the state-of-the-art sequence parallelism methods employed for processing long sequences with transformers.
@@ 4.1 Communication Analysis
The primary advantage of DSP lies in its ability to minimize communication costs and enable scalable communication operations. DSP exploits the inherent characteristics of multi-dimensional transformers to eliminate unnecessary communication, compared with embedded approaches such as Megatron-SP [22], RingAttention [25] and DeepSpeed-Ulysses [20]. Consider an activation size of $M$ and a sequence parallel size of $N$. In a 2D-Transformer, there is one transformer block for each sequence dimension per layer, resulting in two transformer blocks per layer. More details of the implementation of each method are demonstrated in Appendix A.2.
Both Megatron-LM and DeepSpeed-Ulysses require transforming sequence parallelism from a sequence-shard to a head-shard layout. Megatron-SP employs 2 all-gather operations to aggregate the entire sequence and 2 reduce-scatter operations to distribute results in the attention and MLP layers for one transformer block. This results in a total of 8 collective communication operations, leading to a total communication volume of $8M$. Conversely, DeepSpeed-Ulysses incurs 4 communication operations per block for the query, key, value, and output layout transformations, totaling 8 communication operations. Consequently, the communication volume transmitted per device for an all-to-all communication of size $M$ across $N$ GPUs is $8M/N$. Ring-Attention needs to communicate the entire key and value in the attention that sequence dimension is shard, resulting in a total communication volume of $2M$.
In contrast, DSP mitigates communication cost by employing only two all-to-all operations in total two blocks per layer. As shown in Table 3, it reduces the communication volume to $2M/N$, significantly outperforming other sequence parallelism techniques. Notably, with merely two all-to-all operations, DSP exhibits efficient scalability even in super-large clusters for training and inference on extremely long sequences because the communication volume decreases as the number of nodes increases, rendering DSP an exceptional choice for large-scale distributed training and inference tasks involving extreme long sequences.
Table 3: Comparison of DSP with other sequence parallelism methods for 2D-Transformer architectures. $M$ denotes the activation size, and $N$ represents the number of devices. Communication volume refers to the per-layer volume.
@@ 4.2 Memory Analysis
Regarding activation memory, since we shard every tensor in the transformer, we are theoretically able to achieve the minimum activation cost, similar to DeepSpeed-Ulysses and Ring Attention. In practice, however, our approach requires less shape transformation and communication overhead, allowing us to further reduce intermediate activation memory compared to other methods. Megatron-SP, on the other hand, needs to hold the entire activation after the all-gather operation, resulting in higher memory requirements for hosting the activation.
As for parameter memory, as discussed in Section 3.4, DSP is compatible with most parameter parallelism strategies. In this analysis, we utilize the ZeRO technique [36] to evenly shard all parameters across different devices. Consequently, our parameter memory footprint can be kept low.
@@ 5 Experiments
Experiments are conducted on 128 NVIDIA H100 GPUs, interconnected via NVLink within nodes and InfiniBand across nodes. Our methods and implementations are not dependent on specific hardware architectures and can generalize to other devices, particularly those with less efficient interconnects. We follow the same baseline and settings as discussed in Section 4, utilizing 720M and 3B size Transformer-2D models in our experiments. Despite the existence of various 2D-Transformer variants, their architectures are fundamentally similar. Consequently, we select one base model similar to OpenSora [50], an open-source video generation model, for our study. The code is implemented using PyTorch [33].
@@ 5.1 End-to-End Performance
Figure 5: End-to-end performance comparison of different sequence parallel methods combined with data parallel on 128 NVIDIA H100 GPUs. The sequence parallel size is set to minimum for different method.
@@ 5.2 Scaling Ability
Weak Scaling. In the weak scaling experiments, to maintain a consistent computational workload for each GPU, the batch size is linearly increased proportional to the number of GPUs, while the sequence length is fixed. As shown in Figures 6, DSP significantly outperforms other methods by more than 80.7%. Moreover, DSP can scale up to 64 GPUs without being limited by the number of attention heads, unlike DeepSpeed-Ulysses and Megatron-SP. Despite scaling to 64 GPUs, DSP maintains an almost linear throughput increase, with only a 15% performance loss from 8 GPUs to 64 GPUs. Additionally, DSP can achieve super-linear scaling for intra-node setups due to reduced optimizer and gradient computation overhead, and its highly efficient performance within a single node.
Strong Scaling. In the strong scaling experiments, both the batch size and sequence length are fixed. As shown in Figure 7, DSP can maintain linear scalability when scaling up to 8 GPUs for the 720 M model and 4 GPUs for the 3B model, which covers most practical scenarios. To evaluate the extreme performance capabilities of DSP, we further scale up to 64 GPUs with very little workload per device. Although there is an inevitable performance drop, DSP’s throughput remains significantly better than the baselines, outperforming Ulysses by up to 62%. As shown in Figure 9, we also compare the inference latency of all baselines with the same workload. DSP can speed up 29% to 63% compared with baselines.
@@ 5.3 Memory Consumption
Figure 9 demonstrates the memory consumption comparison of different baselines in the weak scaling setting. The semi-transparent bar represents the cache memory, while the solid bar represents the allocated memory. The total memory usage is the sum of them. Our approach exhibits the lowest memory usage, scaling efficiently for longer sequences. Furthermore, DSP’s memory usage is compact without excessive cache memory bloat, unlike Ring-Attention and Megatron-SP.
6 Conclusion and Discussion
In this work, we introduced Dynamic Sequence Parallelism (DSP), a novel sequence parallel abstraction for effectively scaling multi-dimensional transformers to long sequences. Unlike current embedded sequence parallelism methods that only shard on single sequence dimension and are tailored to specific tasks and patterns, DSP offers a general and elegant solution by dynamically switching the parallel dimension during computation, using an efficient resharding strategy decoupled from the computation module. The key advantages of DSP are: 1) substantially reduced communication costs, 2) adaptability across modules without specialized modifications, and 3) remarkable ease of implementation enabled by a simple high-level API. Our experiments demonstrated DSP’s superiority, achieving from 32.2% to 10x higher end-to-end throughput and at least 75% lower communication volume compared to state-of-the-art methods. Its elegance and ease of use make it a promising solution for efficient sequence parallelism across a wide range of applications.
@@ 6 Conclusion and Discussion
Acknowledgements
Yang You’s research group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, ARCTIC grant, SMI grant (WBS number: A-8001104-00-00), Alibaba grant, and Google grant for TPU usage.
@@ A.1 Model Details
In the theoretical analyses and evaluation section, we use a Transformer-2D model as our base model, similar to OpenSora [50]. However, it is not exactly OpenSora; we have removed its specific cross-attention module to ensure that the performance can be generalized to other models. Therefore, in each layer, there are only two transformer blocks that process two sequence dimensions separately, as shown in Figure 10. Specifically, the two dimensions are temporal $t$ and spatial $s$ for a sequence. Each dimension is processed by a corresponding transformer block, which is a common strategy in many applications.
A.2 Parallelism Implementation
In Figure 10, we demonstrate the detailed implementation of different sequence parallel methods on 2D-Transformer. The implementation of DeepSpeed-Ulysses [39] is directly adopted from the official OpenSora [50] implementation, while Megatron-SP [22] is adopted based on its official implementation. For Ring-Attention [25], we adopt an unofficial implementation for the 2D-Transformer.
Megatron-SP employs four resource-intensive collective communication operations per transformer block. Specifically, it initiates an AllGather operation to aggregate the entire input $x$, succeeded by ReduceScatter operations at the output for both attention and MLP modules, culminating in a total communication volume of $8M$.
DeepSpeed-Ulysses adopts the more efficient AlltoAll approach, augmenting the number of communication operations from two to four per block. It leverages all-to-all for query, key, value to transform their shard dimension before attention, and a all-to-all for output after attention. Consequently, the communication volume transmitted per device for an AlltoAll communication of size $M$ across $N$ GPUs is $8M/N$.
A.3 Experiment Settings
A.3.1 Model Size
Table 4: Model settings of 720M and 3B 2D-Transformer.
Model Name Layers Hidden States Attention Heads Patch Size
720M 28 1152 16 (1, 2, 2)
3B 36 2038 32 (1, 2, 2)
A.3.2 End-to-end Performance
@@
A.3.3 Scaling Ability
@@
A.4 API
Figure 11: An example to demonstrate DSP’s API for Pytorch. Developers can use only four functions to enable DSP. $shard$ refers to the sequence parallel dimension.
In the API, as illustrated in Figure 11, we provide four functions for users: $dsp\_dataloader$, $split$, $gather$, and $dynamic\_switch$. The $dynamic\_switch$ function will reshard sequences between the computation stages. Users only need to input the source shard dimension and target shard dimension. The $split$ and $gather$ functions assist users in distributing and retrieving the entire sequence when needed. The $dsp\_dataloader$ function helps users process their dataloader for sequence parallelism, which requires the same data within a single sequence parallel group.

### v5

@@ 4 Theoretical Analysis
We choose 2D-Transformer as described in Equation 8 as our base model, which is widely employed in real-world applications. To be specific, we use the OpenSora (Zangwei Zheng, 2024) variant of 2D-Transformer, an open-source video generation model, where there are two transformer blocks for two sequence dimensions separately. More details can be found in Appendix A.1. We choose state-of-the-art sequence parallel methods including DeepSpeed-Ulysses (Jacobs et al., 2023), Megatron-SP (Korthikanti et al., 2022), Megatron-LM (Shoeybi et al., 2019) and RingAttention (Liu et al., 2023a) as baselines.
@@ 4.1 Communication Analysis
Table 3: Comparison of DSP with other sequence parallelism methods for 2D-Transformer architectures. $M$ denotes the activation size, and $N$ represents the number of devices. Communication volume refers to the per-layer (two 1D blocks) volume per device.
@@
[…]
The primary advantage of DSP lies in its ability to minimize communication costs and enable scalable communication operations. DSP exploits the inherent characteristics of multi-dimensional transformers to eliminate unnecessary communication, compared with embedded approaches such as Megatron-LM (Shoeybi et al., 2019), Megatron-SP (Korthikanti et al., 2022), RingAttention (Liu et al., 2023a) and DeepSpeed-Ulysses (Jacobs et al., 2023). Consider an activation size of $M$ and a sequence parallel size of $N$. In a 2D-Transformer, there is one transformer block for each sequence dimension per layer, resulting in two transformer blocks per layer (two 1D blocks). More details are demonstrated in Appendix A.3.
Megatron-LM & Megaton-SP
both need to gather and scatter the whole sequence. Megatron-LM utilizes 2 all-reduce per block for attention and mlp, while Megatron-SP employs 2 all-gather and 2 reduce-scatter operations per block. They lead to a total per-device communication volume of $8M$ for both methods.
Ring-Attention
needs to communicate the entire key and value in the temporal block only, resulting in a total per-device communication volume of $2M$.
DeepSpeed-Ulysses
incurs 4 all-to-all in temporal block for the query, key, value, and output of attention, resulting in a total per-device communication volume of $4M/N$.
DSP
mitigates communication cost by employing only two all-to-all operations in total two blocks per layer. As shown in Table 3, it reduces the communication volume to $2M/N$, significantly outperforming other sequence parallelism techniques. Notably, with merely two all-to-all operations, DSP exhibits efficient scalability even in super-large clusters for training and inference on extremely long sequences because the communication volume decreases as the number of nodes increases, rendering DSP an exceptional choice for large-scale distributed training and inference tasks involving extreme long sequences.
@@ 4.2 Memory Analysis
Regarding activation memory, since we shard every tensor in the transformer, we are theoretically able to achieve the minimum activation cost, similar to DeepSpeed-Ulysses and Ring Attention. In practice, however, our approach requires less reshape and communication overhead, allowing us to further reduce intermediate activation memory compared to other methods. Megatron-SP and Megatron-LM, on the other hand, needs to hold some entire sequences, resulting in higher memory requirements.
As for parameter memory, as discussed in Section 3.4, we utilize the ZeRO technique (Rajbhandari et al., 2019) to shard all parameters across different devices to ensure low parameter memory footprint.
@@ 5 Experiments
Figure 5: End-to-end performance comparison of different sequence parallel methods combined with data parallel on 128 H100 GPUs. The sequence parallel size is set to minimum for each method.
Experiments are conducted on 128 NVIDIA H100 GPUs, interconnected via NVLink within nodes and InfiniBand across nodes. Our methods and implementations are not dependent on specific hardware architectures and can generalize to other devices, particularly those with less efficient interconnects. We follow the same baseline and settings as discussed in Section 4, utilizing 720M and 3B size Transformer-2D models in our experiments. Despite the existence of various 2D-Transformer variants, their architectures are fundamentally similar. We select one model similar to OpenSora (Zangwei Zheng, 2024). The code is implemented using PyTorch (Paszke et al., 2019).
@@ 5.2 Scaling Ability
Weak Scaling. In the weak scaling experiments, to maintain a consistent computational workload for each GPU, the batch size is linearly increased proportional to the number of GPUs, while the sequence length is fixed. As shown in Figures 6, DSP significantly outperforms other methods by more than 80.7%. Moreover, DSP can scale up to 64 GPUs without being limited by the number of attention heads, unlike DeepSpeed-Ulysses and Megatron-SP. Despite scaling to 64 GPUs, DSP maintains an almost linear throughput increase, with only a 15% performance loss from 8 GPUs to 64 GPUs. Additionally, DSP can achieve super-linear scaling for intra-node due to efficient communication.
Strong Scaling. In the strong scaling experiments, both batch size and sequence length are fixed. As shown in Figure 7, DSP can maintain linear scalability when scaling up to 8 GPUs for 720M model and 4 GPUs for 3B model, which covers most practical scenarios. To evaluate the extreme performance capabilities of DSP, we further scale up to 64 GPUs with very little workload per device. Although there is an inevitable performance drop, DSP’s throughput remains significantly better than the baselines. As shown in Figure 9, our work can significantly reduce inference latency compared with baselines with the same workload.
5.3 Memory Consumption
Figure 9 demonstrates the memory consumption comparison of different baselines in the weak scaling setting. The semi-transparent bar represents the cached memory, while the solid bar represents the allocated memory. The total memory usage is the sum of them. Our approach exhibits the lowest memory usage, scaling efficiently for longer sequences. Furthermore, DSP’s memory usage is compact without excessive cache memory bloat, unlike Ring-Attention, Megatron-LM and Megatron-SP.
6 Conclusion
In this work, we introduce Dynamic Sequence Parallelism (DSP), a novel sequence parallel abstraction for effectively scaling multi-dimensional transformers to long sequences. Unlike current embedded sequence parallel methods that only shard on single sequence dimension and are tailored to specific patterns, DSP offers a general and elegant solution by dynamically switching the parallel dimension during computation, decoupled from the computation module.
The key advantages of DSP are: 1) substantially reduced communication costs, 2) adaptability across modules without specialized modifications, and 3) remarkable ease of implementation enabled by a simple high-level API. Our experiments demonstrated DSP’s superiority, achieving from 32.2% to 10$\times$ higher end-to-end throughput and at least 75% lower communication volume compared to state-of-the-art methods. Its elegance and ease of use make it a promising solution for efficient sequence parallelism across a wide range of applications.
Limitations.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
We organize our appendix as follows:
  • Section A.1: Model details.
  • Section A.2: Experiment Settings.
    • Section A.2.1: Model size.
    • Section A.2.2: End-to-end performance.
    • Section A.2.3: Scaling ability.
  • Section A.3: Parallelism implementation.
@@ A.1 Model Details
In the theoretical analyses and evaluation section, we use a Transformer-2D model as our base model, similar to OpenSora (Zangwei Zheng, 2024). However, it is not exactly OpenSora; we have removed its specific cross-attention module to ensure that the performance can be generalized to other models. Therefore, in each layer, there are only two transformer blocks that process two sequence dimensions separately, as shown in Figure 10. Specifically, the two dimensions are temporal $t$ and spatial $s$ for a sequence. Each dimension is processed by a corresponding transformer block, which is a common strategy in many applications.
A.2 Experiment Settings
A.2.1 Model Size
@@
Table 5: Model settings of 720M and 3B 2D-Transformer.
Model Name Layers Hidden States Attention Heads Patch Size
720M 28 1152 16 (1, 2, 2)
3B 36 2038 32 (1, 2, 2)
A.2.2 End-to-end Performance
A.2.3 Scaling Ability
@@
A.3 Parallelism Implementation
In Figure 10, we demonstrate the detailed implementation of different sequence parallel methods on 2D-Transformer. The implementation of DeepSpeed-Ulysses (Rasley et al., 2020), Megatron-SP (Korthikanti et al., 2022) and Megatron-LM (Shoeybi et al., 2019) are adopted based on their official implementation. For Ring-Attention (Liu et al., 2023a), we adopt an unofficial implementation and adapt it to the 2D-Transformer.
Communication Type Communication Volume
all-reduce 2M
all-gather M
reduce-scatter M
all-to-all M/N
Table 8: Per-device communication volume for each operation.
Method Communication Type Communication Times Per Layer (Two 1D Blocks) Time (ms)
DSP all-to-all 2 914
DS-Ulysses all-to-all 4 5828
Megatron-LM all-reduce 4 8412
Megatron-SP all-gather + reduce-scatter 4 51908
Table 9: Comparison of communication methods and performance with a 8M length of sequence.
Megatron-SP employs four resource-intensive collective communication operations per transformer block. Specifically, it initiates an all-gather operation to aggregate the entire input $x$, succeeded by reduce-scatter operations at the output for both attention and MLP modules, culminating in a total communication volume of $8M$ for one layer (two 1D blocks). Note that the communication volume is calculated per device.
Similarly, Megatron-LM employs 2 all-reduce per transformer block, culminating 4 all-reduces, in a total communication volume of $8M$.
DeepSpeed-Ulysses adopts the more efficient all-to-all approach. It leverages all-to-all for query, key, value to transform their shard dimension before attention, and a all-to-all for output after attention. And it only need to communicate in temporal transformer block. Consequently, the communication volume transmitted per device for an AlltoAll communication of size $M$ across $N$ GPUs is $4M/N$.
As shown in Table 8 and 9, we demonstrate the communication volume for each communication method, and their latency for a sequence of 8M length.